# Experimental Analysis of Efficiency of the Messaging Layer Security for Multiple Delivery Services

### v1

@@ 4.2 Measurements
We will measure the following values in our executions:
@@
In the context of this work, we define an user’s latency with respect to a commit as ”the amount of time between the generation of the commit and the moment in which its changes are applied to the user”. Since an MLS group is composed of multiple users, we will measure the average latency and max latency for each commit generated. The latter represents the delay between the generation of a commit and the start of the new epoch for all members.
The time between the generation of a message and its processing is relevant for all messaging applications, but is particularly relevant for handshake messages in MLS: when a member issues a commit, the group remains in an unstable state until all parties have processed the change and thus cannot adequately exchange any other handshake or application messages.
Commit generation and processing times
CGKA protocols that employ tree structures, such as MLS, are claimed to have logarithmic complexity in relation to the number of members [6, 20]. These claims refer to the Communication cost, that is, the amount of information that needs to be exchanged in order to heal the group. This metric’s relation to computational costs is unclear, and thus in this work we analyse both generation and processing time. We will use CPU time to measure both values.
Besides group size, we measure how the choice of paradigm and state of the Ratchet Tree affect generation and processing times.
Message size
Our analysis includes handshake messages - both commits and proposals, using the Average Update cost metric - as well as other message types like Welcome and GroupInfo packages. The latter two include the full Ratchet Tree, so their size is expected to scale poorly as the group grows.
This experimental measurement complements the communication cost that has been widely analysed in the literature. Our measurements can provide specific values that can be fed into communication cost calculations such that the obtained results can be applied to real-world scenarios.
5 Implementation
In this Section we will introduce our implementation of the different Delivery Services we will employ in our execution environment. We have developed our implementation in Rust, using the OpenMLS project [29] as a baseline and incrementally expanding upon it.
We employ two different Delivery Services, each of them with their corresponding Directory. The first employs a centralised server and forwards messages through the MQTT protocol while the second distributes them through a peer-to-peer network. Their implementations are analysed in subsequent sections. Both are based in a Publish-Subscribe architecture.
Figure 3 shows the handling of topics of a Publish/Subscribe Delivery Service for MLS. Every group has a different topic to which all current members are subscribed. Every user is subscribed to a welcome topic through which it receives all Welcome messages directed to them. Additionally, all groups have their own topic that all members are subscribed to, and every handshake and application message of that group is published through it.
Figure 3: Subscription process for new group members.
5.1 MQTT Delivery Service
@@
MQTT is a protocol for establishing publish/subscribe queues of messages [30]. In this architecture, users can publish messages to different topics such that they are received all users that are subscriber to said topic. The transmission of these messages to all subscribers is managed by a centralised server named MQTT broker. For our execution environment, we employ the open-source Mosquitto [31] implementation of a MQTT broker.
@@
Besides the MQTT broker, the Delivery Service also employs a web server that acts as Directory, with endpoints to store and consume KeyPackages and GroupInfo messages. The latter needs to be updated for each new group epoch, so users aiming to perform an External Join can successfully access the group. Furthermore, this web server also acts as a signaling server in which users can register to notify other participants of their existence, such that they can be invited to MLS groups through an Add proposal.
5.2 GossipSub Delivery Service
For this Delivery Service we maintain the Publish-Subscribe nature but replace the centralised MQTT server with a distributed architecture. We employ the GossipSub protocol [32] that allows peers to subscribe to different topics; peers share with each other metadata of messages they have received and can request the full message if it is from a topic they are subscribed to. Connections are dynamic and favour other members subscribed to the same topic.
We also employ a Decentralised Hash Table (DHT) as this Delivery Service’s Directory. Each client publishes its KeyPackage as a key-value pair, with the key referencing its user identifier. We limit the involvement of the server to that of signaling server as defined for the MQTT Delivery Service.
For our implementation we include Rust’s Libp2p [33] into the simulated clients, which provides an implementation for GossipSub and the Kademlia [34] DHT.
5.3 Client Simulation
@@
Our testbed is composed of multiple simulated clients that independently participate in MLS groups. They emulate the behaviour of real clients by randomly adding or removing members, updating their individual leaf nodes and sending application messages.
[cgka] ds = "mqtt" groups = ["group_1", "group_2"] external_join = true join_chance = 0.01 issue_update_chance = 0.2 message_chance = 0.3 scale = false auth_policy = "Random" message_length_min = 500 message_length_max = 2000 sleep_millis_min = 20000 sleep_millis_max = 60000 [paradigm] paradigm = "propose" proposals_per_commit = 4 invite_chance = 0.6 remove_chance = 0.1 update_chance = 0.3 [http_server] url = "http://<ip>:<port>" [mqtt] url = "tcp://<ip>:<port>" [meta] replicas = 10
Figure 4: Example Client configuration file.
The behaviour of our simulated clients can be specified in detail through configuration files, as shown in Figure 4. Among other parameters, it is possible to configure the groups the user will try to join (or create if they do not exist), how often it will attempt to perform an action and with what probability. The configuration file also defines which paradigm to use and the types of actions (Add, Remove, Update, Join) the client is execute, as well as the Delivery Service that will be employed.
group_1 8 User_0d4341e0c0f4 Invite User_b3f20ed42c2a 1065 1739176120380282661 5885 group_1 8 User_d13041578e84 Process User_0d4341e0c0f4 1739176120384444036 8599 group_1 8 User_b3f20ed42c2a Welcome User_0d4341e0c0f4 2755 1739176120461981624 1823
Figure 5: Logs generated by clients.
Every action related to the evolution of an MLS group is recorded in a log file. Figure 5 shows the trace of an Invitation, which includes involves the following entries: Invite, Process and Welcome. Each entry records the size of the message, the timestamp at which it was performed and its time cost.
@@ 6 Experimental Results
In this section we deploy the environment presented in previous sections and perform empirical evaluations on its performance.
We have executed all following tests three times with the objective of clearing any potential noise in the measurements. Every figure in this Section shows the average of the results obtained in all executions; we omit the standard deviation as it is not relevant to our scenarios.
6.1 Latency
(a) MQTT
(b) GossipSub
Figure 6: Mean and Max Latency for the Delivery Services.
Figures 6(a) and 6(b) show the latency for the two Delivery Services. While both experience a significant increase as the number of users grow, the GossipSub DS is noticeably less efficient as its latency more than doubles that of the MQTT DS. Latency also increases more irregularly of GossipSub, with an important increase at 512 group members. However, at higher group sizes the latency grows slower for GossipSub.
6.2 Paradigm
We now compare the Average Update cost for the commit and propose-and-commit paradigms, including multiple amounts of proposals per commit for the latter to see how the cost varies as the number of updates increases. We note that our measurements do not include the cost of retrieving information from the Directory such as KeyPackages for Add proposals, as it remains constant regardless of the chosen paradigm.
(a) Commit and propose-and-commit paradigms
(b) Commits (also including Welcome processing time) and External Joins.
Figure 7: Average Update generation time for different paradigms as the number of users grow.
@@ 6.2 Paradigm
Figure 8: Average Update cost for different paradigms as the number of users grow.
Figures 7(a) and 7(b) show the Average Update generation time for the two paradigms and for External Joins, respectively. Clearly, the Average Update cost is significantly reduced as the number of proposals increases.
We also note that the time to perform an External Join is much larger to that of inviting new members through Add proposals. In order to understand these differences we highlight that an user performing an External Join must also parse the current state of the group including the full Ratchet Tree, which would explain the poor scaling of External Joins. This operation is also performed in an Invitation setting by the invited user when the Welcome message is processed; when accounting for the processing time of Welcome messages, the two methods have a similar cost (as shown in Figure 7(b) by the line labelled Commit + Welcome).
Figure 8(a) shows the processing time per Update as the number of users grow. As with generation time, the processing cost is significantly reduced as the number of proposals increases. Processing time is noticeably more stable than generation time, although it experiences significant increases at the powers of 2 as a new layer is added to the Ratchet Tree. It is apparent that both generation and processing times scale mostly linearly with the number of members in the CGKA group, with $R^{2}$ scores over 0.9 when adjusting for linear regression and around 0.6 for logarithmic regression.
@@
Similarly, Figure 8(b) shows the message size per update. Clearly, the correlation between the message size and the number of users is much smaller than in the other metrics. Contrary to the results obtained while measuring processing and generation times, message size is generally bigger as the number of proposals increase. We recall that commit messages also include copies of the proposals it applies; this redundant information increases the cost per update.
6.3 Other Messages
Figure 9: Message size for Welcome messages and GroupInfo packages.
@@ 6.3 Other Messages
We also evaluate the size of the Welcome messages and GroupInfo packages, which take part in Invitations and External Joins, respectively. Figure 9 shows the obtained results. Since both contain a full copy of the Ratchet Tree, they are significantly larger than any handshake message. Expectedly, their size increases linearly with the number of group members.
6.4 Ratchet Tree State
(a) Generation time.
(b) Message size.
Figure 10: Generation time and message size for different states of the group’s Ratchet Tree. Processing time is omitted as it is not influenced by this metric.
Finally, we analyse the impact in performance of the current state if the Ratchet Tree, exemplified by the First, Last and Random scenarios introduced in Section 4.1. Figure 10 shows the generation time and message size - processing time is omitted as no significant difference was found between the three scenarios.
As with previous measurements, generation time also scales linearly in all scenarios - this includes Last which corresponds to the best possible state of the tree. Message size does show a significant correlation, with First increasing much faster than in the other scenarios.
@@ 7 Discussion
Comparison with other works
Protocol Perspective Group Size Time Bandwidth Tree State Paradigm Join DS
[6] CGKA Variant Theoretical N/A
[3, 4] CGKA Theoretical N/A
[11] CGKA Variant Experimental $\approx$ 128
[10] CGKA Variant Experimental $\approx$ 128
[24] MLS Theoretical N/A
Ours MLS Experimental $\approx$ 1000
Table 1: Comparison between works that analyse CGKA performance and ours. The columns indicate if the works consider the following topics. CPU: update generation/processing time. Bandwidth: message size and communication cost. Paradigm: influence of chosen paradigm. Tree State: Influence of the current state of the Ratchet Tree. Join: External Joins. DS: influence of the Delivery Service.
Table 1 shows a comparison between our work and other works that perform an analysis of performance of CGKA. Our experimental approach allows us to measure CPU time in generating and processing messages and the importance of the Delivery Service in their distribution. When compared with other experimental works, ours has the largest group size and also covers less studied topics such as the influence of paradigm and the current state of the Ratchet Tree.
Delivery Service
The increased bandwidth of larger groups has a noticeable impact on latency, as the cost of distributing messages between all members increases. The MQTT DS performs significantly better than the GossipSub DS, but the former does not require a centralised server to distribute the messages and thus is more reliable.
Latency represents the most significant cost in the evolution of MLS groups, as every other generation or processing cost is negligible in comparison. Updates in MLS are necessarily sequential: if two commits are generated from the same epoch, the group enters into an inconsistent state. In order to generate a valid commit, an user must have received and processed all previous valid commits. Thus, latency is a significant bottleneck to group evolution as it represents the minimum amount of time between two valid commits.
Our measurements reach up to 2 seconds of maximum latency in which the group is in an inconsistent state; while for most applications this span is too short to be noticed, it is something that must be taken into account for situations that require strict synchronisation.
Message cost
Generation and processing times are comparable and represent a small cost in the order of milliseconds. Both experience noticeable increases when a new layer is added to the Ratchet Tree, that is, when group size reaches the next power of 2. In practice, both generation and processing times increase linearly with the amount of users in the group, which contradicts the theoretical assumptions that the protocol complexity is logarithmic. Notably, this trend is maintained even in the best-case scenario where the minimum amount of intermediate nodes are blanked. When compared to latency, with generation and processing times represent a negligible cost.
Welcome messages and GroupInfo packages are significantly costlier to process - as shown in Figure 7(b) - than normal commits. Fortunately, they only need to be processed once in order to access a group. This cost difference is due to their size as they both contain the full Ratchet Tree, which needs to be parsed by the receiving party. In contrast, the size of handshake messages is less influenced by group size and is mostly determined by the state of the Ratchet Tree. Generation time is also significantly affected. Even though these two metrics are slightly correlated they present significant differences and thus should be analysed separately. Our results show that groups in which only a small number of users generate updates are less efficient.
Impact of paradigm
@@
The propose-and-commit paradigm also significantly affects performance: the cost of generating and processing proposal messages is negligible compared to the cost of applying an update to the group. As shown, more proposals per update increase the efficiency of the protocol, but we remark that such amount of modifications to a group applied at once would be rare in a real setting, only possible in a highly unstable group. The size of handshake messages is less affected by the amount of group members and the paradigm, as it is mostly determined by the current state of the Ratchet Tree.
External Joins combine the operations of generating a commit and processing a Welcome message, in the sense that the joiner needs to parse the full Ratchet Tree in order to join. When taking these two operations into account, the cost of an Invitation is similar to that of an External Join.

### v2

@@ 4.3 Measurements
In this work we will focus in computational costs, that is, the time it takes to compute certain MLS operations. In particular, we will obtain generation and processing times for different types of MLS messages, measured in CPU time. Furthermore, we will also analyse the size of MLS message: this includes handshake messages — both commits and proposals — as well as other message types like Welcome and GroupInfo packages.
Update Cost
When considering the impact of paradigm, we unify all costs — both time costs and message sizes — to account for the different amount of proposals being applied in a single commit. Recall that whereas in the commit paradigm every handshake message applies exactly one modification, in the propose-and-commit paradigm a commit that applies $n$ proposals requires the exchange of $n+1$ messages: each of the $n$ proposal messages and the commit message itself. Thus, we define the update cost as
$\mathsf{UC}=\frac{cc+\sum_{i=0}^{n}cp_{i}}{n},$
where $cc$ and $cp_{i}$ are the costs of the commit and its $i$th proposal, respectively. Note that in the commit paradigm, $\sum_{i=0}^{n}cp_{i}=0$ and $n=1$, thus the update cost equals $cc$.
Deep Analysis
The experiments First and Last respectively represent the least and most optimal states of the Ratchet Tree. As shown in Figure 4, this impacts the amount of ciphertexts that must be created to communicate the new state of the group. In order to further understand the differences between both experiments, we will additionally perform a deeper analysis for these experiments that subdivides the measurements obtained into categories, each representing a different aspect of the creation of a commit:
  • 1.
    Validation: Operations to ensure that the proposals to apply are valid. Of particular interest is the MLS requirement that encryption keys be unique among all members.
  • 2.
    Cryptography: TreeKEM operations such as updating and encrypting the intermediate nodes’ secrets and updating the key schedule plus signing and encrypting MLS messages.
  • 3.
    Tree: Operations related to parsing the Ratchet Tree such as calculating the path to the root and finding which keys to use to encrypt and decrypt path secrets. Additionally, operations that maintain the integrity of the tree, such as and updating the tree hashes and parent hashes every epoch [12].
  • 4.
    Welcome: Operations to communicate the state of the group to a new member, such as exporting the GroupInfo package and the Ratchet Tree. Only applicable to Commit generation.
  • 5.
    Storage: Operations to provide persistence to the application by saving the group’s state into memory.
@@
Finally, we also measure latency, defined as the time between the instant a message is sent by a committer and the instant it is received by a processer. Whenever a commit is issued but has not been processed yet by all members, the group is in an inconsistent state as members disagree on the current epoch and group state: latency determines how long this inconsistency may last. To avoid redundancy, latency does not include neither commit generation nor processing times.
5 Emulation Environment
In this Section we introduce the implementation of the presented testbed. Our code [^2] Available at https://github.com/SDABIS/mls_experimental_analysis consists mostly of Rust code that implements the emulated MLS client, as well as the Docker infrastructure required to execute the Delivery Service and multiple clients.
5.1 Delivery Service
@@
MQTT is a protocol for establishing publish/subscribe queues of messages [28]. The transmission of these messages to all subscribers is managed by a MQTT broker. For our execution environment, we employ the open-source Mosquitto [27] implementation of a MQTT broker.
@@
Besides the MQTT broker, the Delivery Service also employs a web server that acts as Directory, with endpoints to store and consume KeyPackages and GroupInfo messages. The latter needs to be updated for each new group epoch, so users aiming to perform an External Join can successfully access the group. Furthermore, this web server also acts as a signaling server in which users can register to notify other participants of their existence, such that they can be invited to MLS groups.
5.2 Client Emulation
@@
Our testbed is composed of multiple emulated clients that independently participate in MLS groups. They emulate the behaviour of real clients by randomly adding or removing members, updating their individual leaf nodes and sending application messages. Clients interact with each other through real MLS messages sent through the Delivery Service.
We have developed our implementation of an emulated client in Rust, using the OpenMLS project [31] as a baseline and incrementally expanding upon it. Clients can be deployed in Docker containers, each executing a configurable amount of copies. The behaviour of Clients can be extensively configured to reach specific group states. Each client has a different name and a list of groups they aim to join.
There are two behaviours clients can follow to decide when to act in a group. In the independent mode, clients randomly decide whether or not to send a message at fixed time intervals. This may result in redundant messages if multiple users attempt to commit at the same time. In orchestrated mode, the transcript hash of the group is used as a seed to a RNG, such that all group members agree on a randomly chosen client who issues the next update. This allows for better coordination resulting in faster executions. In either mode, clients always process incoming messages immediately after their arrival. Clients can also be configured to only consider issuing updates if they are the first or last member to join the group, as required by the Tree Degradation experiments described in Section 4.
When they are tasked with issuing updates, clients will follow one of the paradigms described in Section 4. If the propose-and-commit paradigm is employed, clients will issue proposals until a threshold of proposals is reached and then commit them. Clients may also perform an External Join to a group in which they are not members.
Clients randomly choose between adds, updates and removals according to configured proportions. In order to reach groups of stable size for the experiments introduced in Section 4, this proportions can be overridden once the group reaches a threshold of members such that only removes may be issued. Additionally, clients may send randomly-generated application messages. The emulated clients record every action they perform in detailed log file that includes the timestamp and duration of the action.
@@ 6 Experimental Results
In this section we deploy the environment presented in previous sections and perform empirical evaluations on its performance. All experiments in Table1 have been executed by providing their respective configuration to the emulated clients. The target group size for all experiments —except those related to stable group size— is 10000. We note that in all experiments clients employ the orchestrated behaviour since it allows for faster and more controlled executions.
All experiments were executed on an Ubuntu 20.04 virtual machine equipped with 128 GB of RAM and 16 CPU cores operating at 3.2 GHz. The 10,000 emulated clients were divided into 10 Docker containers each running 1000 instances. The Delivery Service —including both the HTTP Directory and the MQTT Broker— were run in a similar virtual machine located in the same local network.
6.1 Group Evolution
In order to ensure groups grew in size, clients were configured with probabilities of 60% invite, 20% update and 20% remove. Every experiment has been executed 3 times; the data shown in this Section are the aggregation of the results obtained in every execution. For each of the following graphs, the 10,000 measurement points have been subsequently aggregated into 100 points, with each point representing the aggregation of 100 measurements.
Figure 5 shows the mean and maximum latency in the experiment Commit. Both measurements increase slowly as the group grows and the consumed bandwidth increases.
Figure 5: Mean and Maximum Latency for the Delivery Service as the number of group members grows.
6.1.1 Paradigm
We now compare the update cost in our Paradigm experiments. We note that our measurements do not include the cost of retrieving information from the Directory such as KeyPackages for Add proposals, as it remains constant regardless of the chosen paradigm.
(a) Experiments Commit, 2 Prop, 4 Prop and 8 Prop.
(b) Experiment External Join, with Commit shown for comparison. Welcome also including Welcome processing time — and External Joins.
Figure 6: Update generation time for different paradigms as the number of group members grow. Update cost is calculated by dividing total cost by number of proposals applied in a single commit.
Figures 6(a) shows the generation time for the experiments Commit, 2 Prop, 4 Prop and 8 Prop, whereas Figure 6(b) shows the same measure for the experiment External Join. Clearly, the update cost is significantly reduced as the number of proposals increases. We highlight that an user performing an External Join must also parse the current state of the group including the full Ratchet Tree, which would explain the poor scaling of External Joins. This operation is also performed in an Invitation setting by the invited user when the Welcome message is processed; when accounting for the processing time of Welcome messages, the two methods have a similar cost (as shown in Figure 6(b) by the line labelled Commit + Welcome).
@@ 6.1.1 Paradigm
Figure 7: Update cost for different paradigms as the number of users grow. Update cost is calculated by dividing total cost by number of updates of proposals applied in a single commit.
@@
Figure 7(a) shows the processing time per Update as the number of users grow. As with generation time, the processing cost is significantly reduced as the number of proposals increases. Similarly, Figure 7(b) shows the message size per update. Clearly, the correlation between the message size and the number of users is much smaller than in the other metrics. The increases at powers of 2 is also noticeable in message size, as shown by the spikes at 4096 and 8192 members.
@@
We also evaluate the size of the Welcome messages and GroupInfo packages, which take part in Invitations and External Joins, respectively. Figure 8 shows the obtained results. Since both contain a full copy of the Ratchet Tree, they are significantly larger than any handshake message.
Figure 8: Message size in Bytes for Welcome messages and GroupInfo packages.
6.1.2 Tree Degradation
Figure 9: Message size for the First and Last experiments as the number of users grow. The results of the Commit experiment are included for comparison.
Figure 9 shows the message size for the experiments First and Last. Since in the former almost all intermediate nodes are blanked, the commit messages must contain one ciphertext for every member and thus their size grows linearly.
(a) Experiment First.
(b) Experiment Last.
Figure 10: Generation time measured for the experiments First and Last that respectively correspond to the worst and best possible Ratchet Tree states. Measures are subdivided into categories. Storage is cropped to improve visibility, as it increases linearly and experiences significant jumps at powers of two.
Figure 10 shows the distribution of time cost across different categories for commit generation in both experiments. Whereas most categories predictably exhibit the same behaviour, Cryptography is significantly more costly in First than in Last, as the path secrets need to be encrypted once for every member. In both, Processing time is omitted as the cost distribution across categories was identical to that of commit generation. Notably, Cryptography represents a small percentage of the total cost: at 10,000 users, it reaches 1% for the experiment Last and 10% for First.
6.2 Stable group size
In this Section we execute the experiments related to stable group size: 100 Users, 750 Users and 5000 Users. For each experiment we measure 10,000 commits at the corresponding group size. Figure 11 shows the measured commit generation and processing time and message size. The latter metric experiences a noticeably higher variance than the others. In particular, commit processing is the more stable metric. All of them follow a normal distribution, which is more easily observable for larger groups.
(a) Commit size.
(b) Generation time.
(c) Processing time.
Figure 11: Commit size and generation and processing time for groups of 100 (top), 750 (middle) and 5000 (bottom) members. The dotted and dashed lines respectively represent the median and mean of each series.
@@ 7 Discussion
In this Section we analyse the data introduced in the previous Section and discuss its real-world implications. We also compare our work to other analysis of performance of MLS or other CGKA variants. Finally, we discuss the limitations encountered and propose future work.
7.1 Analysis
@@
Message cost. Generation and processing times are comparable and represent a small cost in the order of milliseconds. The propose-and-commit paradigm also significantly affects performance: the cost of generating and processing proposal messages is negligible compared to the cost of applying an update to the group. As shown, more proposals per update increase the efficiency of the protocol. The size of handshake messages is less affected by the amount of group members and the paradigm, as it is mostly determined by the current state of the Ratchet Tree. External Joins combine the operations of generating a commit and processing a Welcome message, as the joiner needs to parse the full Ratchet Tree in order to join. When taking these two operations into account, the cost of an Invitation is similar to that of an External Join.
As shown in Figure 11, commit messages at a given group size can have significantly different sizes. This is determined by the number of ciphertexts contained in them, which in turn depends on the number of blanked intermediate nodes as explained in Figure 4. Notably, commit generation and processing times have less variance as their cost is mostly determined by other operations unrelated to the cryptographic component of MLS.
The increased bandwidth of larger groups has a noticeable impact on latency, as the cost of distributing messages between all members increases. However, this metric grows slower than commit generation and processing times. When comparing Figure 5 with Figures 6 and 7, mean latency is around 2.5 times more expensive than commit generation and processing at 1000 members, whereas their cost is comparable at 10,000 members.
Ratchet Tree Size. Our results show that the overwhelming majority of the costs both in generation and processing are caused by the size of the full Ratchet Tree. Indeed, Figure 10’s categories Storage, Welcome, Validation and Tree all scale linearly with the size of the group’s state; only Cryptography is independent, but its cost is negligible in comparison. The result is that the overall scaling of MLS’s performance is linear, which contradicts the theoretical results which point to logarithmic performance [10, 9, 16]. We remark that Storage cannot be considered optional in practical scenarios, as it is required to provide persistence between sessions. Even then, the remaining categories still scale linearly and their cost cannot be circumvented: Tree includes the calculation of tree hashes and parent hashes that prevent insider attacks [8] and Validation’s check that all keys should be different is needed to ensure Post-Compromise Security.
Furthermore, the most expensive operations of the protocol are processing a Welcome and performing an External Commit, as shown in Figure 6(b). Both require processing the full Ratchet Tree, whether it be from a Welcome message or a GroupInfo package. Thus, the most relevant factor to consider regarding performance in MLS is the size of the Ratchet Tree, which scales linearly by the number of members. The logarithmic complexity of the cryptographic operations only applies to a negligible step of the computation.
7.2 Real-world Implications
Figure 5 shows that maximum latency can reach up to 0.3 seconds, which could potentially be much higher depending on the network or Delivery Service employed. This measure determines the amount of time in which the group is in an inconsistent state, since some members have not yet joined the current epoch. Whereas for most applications this span is too short to be noticed, it is something that must be taken into account for situations that require strict synchronisation among group members.
The results of the experiment First show that groups in which only a small number of users generate updates are less efficient, with a particularly high commit size. This is significant as real-world messaging applications usually limit updates to one or more administrators. Thus, members should be allowed to periodically update their own leaf node to prevent the accumulation of blanked intermediate nodes.
As shown in Figure 10, the size of the Ratchet Tree has a significant impact in performance as this structure needs to be managed to maintain its integrity and stored in memory at every commit. This means elements that appear unrelated to performance but increase the size of the tree (e.g. clients using large credentials or signature keys) may provoke an unexpected increase in the cost of group operations.
7.3 Comparison with other works
Table 2 shows a comparison between our work and other works that perform an analysis of performance of CGKA. Our experimental approach allows us to measure CPU time in generating and processing messages. Our work covers topics whose impact in performance has not been studied in any other of the listed works, such as External Joins, the choice of paradigm and the importance of the Delivery Service.
Notably, other works that analyse the theoretical performance of CGKA [10, 9, 4, 17] focus on communication cost, which is defined as the number of ciphertexts required to heal from corruption. Our experimental analysis serves as a complement to these theoretical bounds by providing a real-world reference of the size of the exchanged messages as well as the cost of generating and processing them. Furthermore, a theoretical approach does not adequately capture the cost of maintaining and storing the Ratchet Tree, which our analysis shows to be much higher than that of cryptographic functions.
Work Protocol Perspective Group Size Time Bandwidth Tree State Paradigm Join DS
Alwen et al. [7] Variant (SAIK) Theoretical N/A
Auerbach et al. [10] CGKA Theoretical N/A
Anastos et al. [9] CGKA Theoretical N/A
Weidner et al. [34] Variant (DGKA) Experimental $\approx$ 128
Chevalier et al. [18] MLS Experimental $\approx$ 1,000
Balbás et al. [11] Variant (A-CGKA) Experimental $\approx$ 128
Paillat et al. [30] MLS Theoretical N/A
Ours MLS Experimental $\approx$ 10,000
Table 2: Comparison between works that analyse CGKA performance and ours. The columns indicate if the works consider the following topics. Time: update generation/processing time. Bandwidth: message size and communication cost. Tree State: Influence of the current state of the Ratchet Tree. Paradigm: influence of chosen paradigm. Join: External Joins. DS: influence of the Delivery Service.
When compared with other experimental evaluations such as [34, 18, 11], the increased depth of our analysis allows us to extract more robust conclusions. Indeed, the increments in computing cost at group sizes that are powers of 2 are not captured by works that limit this metric to 100 or 1,000 members. Our analysis also shows that cryptographic operations represent a very small percentage of experimental cost. Thus, claims about a CGKA variant’s efficiency must take into account the disproportionate cost of storage and tree management that may be unrelated to its contribution.
7.4 Limitations
In this work, we employ an ideal Delivery Service that is readily accessible to all members and that ensures that messages are delivered in order. Thus, our analysis does not capture real-world inconveniences that can be introduced by the network such as limited bandwidth, packet losses or out-of-order delivery. MLS has also sparked interest for its application in distributed environments [3, 32, 26]; replacing the centralised Delivery Service employed in our testbed with a distributed one could be of interest to analyse its implications on performance.
Our analysis is limited to 10,000 users because of the significant cost of the execution of the experiments. For reference, a single commit in a 10,000-member group takes 0.1 seconds to be processed: if executed sequentially, it takes more than 15 minutes for all members to process it. In order to measure performance in larger groups, a stronger effort in parallelism is required. Alternatively, larger analysis could also ignore memory storage in order to remove the largest computational cost.
The results obtained in this work show the relevance of handling the state of the Ratchet Tree. Thus, future work could focus on studying alternatives that minimise its size, such as compression. Similarly, the performance of clients that do not possess a copy of the Ratchet Tree —such as the recently proposed Partial MLS [24]— could be compared to our results.